The generating function of the sequence {1,2,3...n..} is (1-z)².
1
Expert's answer
2020-12-11T13:08:54-0500

No.


We know that


"\sum_{n=0}^{\infin} {z^n} =\frac 1 {1-z}"


Find derivative of it:


"\sum_{n=1}^{\infin} {n*z^{n-1}} =\frac 1 {(1-z)^2}"


Change n to n+1


"\sum_{n=0}^{\infin} {(n+1)*z^{n}} =\frac 1 {(1-z)^2}"


Hence, the generating function of the sequence {1,2,3...n..} is "\frac 1 {(1-z)^2}"